SBD Phone Company sells its waterproof phone case for $116 per unit. Fixed costs total $227,000, and variable costs are $58 per unit. 


(1) Determine the contribution margin ratio. 

(2) Determine the break-even point in dollars. 

SBD Phone Company sells its waterproof phone case for $107 per unit. Fixed costs total $197,000, and variable costs are $32 per unit. Compute the units of product that must be sold to earn pretax income of $224,500. 

Units to be sold to achieve targeted income 

Choose Numerator:/Choose Denominator:=Units to achieve target

/
=Units to achieve target




0
Answer #1

1)

(1) Determine the contribution margin ratio
per unit
Sales $116
Less: Variable costs 58
Contribution margin $58
Contribution margin ratio
Choose Numerator: / Choose Denominator: = Contribution margin ratio
Contribution margin / Sales = Contribution margin ratio
$58 / 116 = 50%
(2) Determined the break-even point in dollars.
Choose Numerator / Choose Denominator = Break-even point in dollars
Fixed cost / Contribution margin ratio = Break-even point in dollars
$227000 / 50% = $454000

2)

Units to be sold to achieve targeted income
Choose Numerator: / Choose Denominator = Units to achieve target
Total fixed costs plus pretax income / Contribution margin per unit = Units to achieve target
$421500 / 75 = 5620 units

Total fixed costs plus pretax income= Fixed costs+Pretax income

= $197000+224500= $421500

Contribution margin per unit= Sales price per unit-Variable costs per unit

= $107-32= $75